OpenStudy (raheelafzaal):

derivation of sin^2x?

OpenStudy (istim):

I'm just learning about this too. See if I answer this right. Is it 2xsin^2x-1?

OpenStudy (raheelafzaal):

2sinxcosx am i right?

OpenStudy (anonymous):

sin(A + B) = sin A cos B + cos A sin B sin(2x) = sin(x + x) = sin x cos x + cos x sin x = 2 sin x cos x SOURCE http://answers.yahoo.com/question/index?qid=20081116100909AA0li2n

OpenStudy (istim):

Oh well.

OpenStudy (mimi_x3):

its 2sinxcosx

OpenStudy (raheelafzaal):

thanks

OpenStudy (raheelafzaal):

\[\int\limits_{0}^{pi}cosxdx?\]

OpenStudy (raheelafzaal):

\[\int\limits_{0}^{pi/2}cosxdx?\]

OpenStudy (mimi_x3):

well, when you integrate cosx its sinx. then you sub in the limits.

OpenStudy (anonymous):

first integrate integral of cosx=sinx now put in the limits sin(pie/2) - sin(0)=1 answer :D